# Exploring the quantum capacity of a Gaussian random displacement channel using Gottesman-Kitaev-Preskill codes and maximum likelihood decoding

### v1

@@ Abstract
Introduction–Quantum communication Bennett and Wiesner (1992); Bennett et al. (1993) and quantum computing Shor (1994); Lloyd (1996); Nielsen and Chuang (2000) are two intricately related fields within quantum information science, aiming to leverage the principles of quantum mechanics to revolutionize data transmission and information processing. One common challenge in these areas is the need to protect the encoded quantum information from noise in quantum systems. Quantum error correction (QEC) thus plays a pivotal role in both fields by providing a mean to systematically suppress the effects of noise. Owing to this close relationship, the concepts and methodologies developed in one area can often be applied to another. For example, noisy processes in quantum computers can be modeled as noisy quantum channels for transmitting quantum information. This analogy is particularly fruitful when it is complemented by a general framework of quantum communication theory. In particular, the latter allows one to study maximum achievable data transmission rates of a noisy quantum channel through various entropic quantities Schumacher and Nielsen (1996); Lloyd (1997); Barnum et al. (1998); Devetak (2005); Wilde (2013), providing fundamental bounds on the capability of a QEC code for correcting errors in the noisy quantum channel.
@@
Bosonic Gaussian channels Weedbrook et al. (2012) have been extensively studied in the field of quantum communication theory Holevo and Werner (2001); Gottesman et al. (2001); Harrington and Preskill (2001); Harrington (2004); Caruso and Giovannetti (2006); Wolf et al. (2007); Holevo (2007); Pirandola et al. (2017); Sharma et al. (2018); Rosati et al. (2018); Noh et al. (2019, 2020); Fanizza et al. (2021); Mele et al. (2022, 2024) due to their relevance to realistic optical and microwave quantum communication channels. Moreover dominant errors in bosonic modes in various quantum computing platforms (e.g., 2D resonators Putterman et al. (2024) and 3D cavities Paik et al. (2011); Reagor et al. (2013); Milul et al. (2023) in superconducting systems, motional modes in trapped-ion systems Flühmann et al. (2019); Matsos et al. (2024), and light modes in optical systems Madsen et al. (2022)) are well described by bosonic Gaussian channels. Thus, numerous bosonic QEC codes Cochrane et al. (1999); Jeong and Kim (2002); Gottesman et al. (2001); Leghtas et al. (2013); Michael et al. (2016); Albert et al. (2018) have been proposed to correct for errors in these bosonic Gaussian channels.
@@
In this work, we explore the interface between quantum communication theory and quantum error correction through bosonic Gaussian channels and bosonic QEC. In particular, we consider the problem of determining the quantum capacity Schumacher and Nielsen (1996); Lloyd (1997); Barnum et al. (1998); Devetak (2005) $Q(\mathcal{N}_{\sigma})$ of a Gaussian random displacement channel $\mathcal{N}_{\sigma}$, i.e., its maximum achievable quantum state transmission rate per channel use under an optimal QEC strategy. The quantum capacities of some special cases of bosonic Gaussian channels, such as the pure-loss and pure-amplification channels, are well understood analytically Holevo and Werner (2001); Wolf et al. (2007); Pirandola et al. (2017); Sharma et al. (2018) thanks to their degradability or anti-degradability property Caruso and Giovannetti (2006). However generic bosonic Gaussian channels including the Gaussian random displacement channel do not possess this special property. Thus only lower Holevo and Werner (2001); Noh et al. (2020) and upper Holevo and Werner (2001); Pirandola et al. (2017); Sharma et al. (2018); Rosati et al. (2018); Noh et al. (2019) bounds on their quantum capacity are known.
@@
Here we attempt to make progress on this long-standing open problem by investigating the performance of multi-mode Gottesman-Kitaev-Preskill (GKP) codes Harrington and Preskill (2001); Harrington (2004); Fukui et al. (2017, 2018); Fukui (2023); Vuillot et al. (2019); Noh and Chamberland (2020); Hänggli et al. (2020); Royer et al. (2022); Conrad et al. (2022, 2024); Lin et al. (2023) against errors caused by a Gaussian random displacement channel $\mathcal{N}_{\sigma}$. Specifically, we aim to find a tighter lower bound on the quantum capacity of the Gaussian random displacement channel by searching for explicit QEC schemes which achieve higher quantum state transmission rates than the best-known lower bound on $Q(\mathcal{N}_{\sigma})$. Our search is based on multi-mode GKP codes and their maximum likelihood decoding.
@@
Gaussian random displacement channel and scaled self-dual GKP codes–In the Heisenberg picture, a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ transforms the position and momentum operators $\hat{q}$ and $\hat{p}$ of a bosonic mode into $\hat{q}+\xi_{q}$ and $\hat{p}+\xi_{p}$, respectively. Here, $\xi_{q}$ and $\xi_{p}$ are independent classical noise variables following a Gaussian distribution with vanishing mean and a variance of $\sigma^{2}$. Note that a Gaussian random displacement channel is also referred to as an additive Gaussian noise channel in the literature.
@@
Here, $I_{c}(\mathcal{N},\hat{\sigma})\equiv S(\mathcal{N}(\hat{\sigma}))-S(\mathcal{N}^{c}(\hat{\sigma}))$ is the coherent information of the channel $\mathcal{N}$ with respect to an input state $\hat{\sigma}$, $S(\hat{\sigma})\equiv-\text{Tr}[\hat{\sigma}\log_{2}\hat{\sigma}]$ is the quantum von Neumann entropy, and $\mathcal{N}^{c}$ is the complementary channel of $\mathcal{N}$ Schumacher and Nielsen (1996); Lloyd (1997); Barnum et al. (1998); Devetak (2005). However in practice, evaluating this entropic quantity is difficult due to the maximization over all possible input states $\hat{\rho}$ as well as the need to consider an asymptotic limit of $N\rightarrow\infty$.
@@
where the lower bound is obtained by computing its one-shot coherent information (i.e., a similar expression as in Eq. (1) but with $N=1$) Holevo and Werner (2001) and the upper bound is derived using a data-processing inequality Noh et al. (2019). From the lower bound and the definition of the quantum capacity, it follows that there must exist a QEC scheme that achieves a quantum state transmission rate of $\log_{2}(\frac{1}{e\sigma^{2}})$ for the Gaussian random displacement channel $\mathcal{N}_{\sigma}$ with $\sigma\leq 1/\sqrt{e}\simeq 0.6065$. Earlier works showed that scaled self-dual GKP codes can achieve the rate of $\log_{2}(\lfloor\frac{1}{e\sigma^{2}}\rfloor)$ Harrington and Preskill (2001); Harrington (2004). Notably, the floor function here is due to the fact that scaled self-dual GKP codes are restricted to cases where every single one of the $N$ normal modes encodes the same number (a positive integer $\lambda$) of logical states, thereby encoding $\lambda^{N}$ logical states or $N\log_{2}\lambda$ logical qubits overall. Thus when $\frac{1}{e\sigma^{2}}$ is given by a positive integer (e.g., $\frac{1}{e\sigma^{2}}=2\leftrightarrow\sigma=1/\sqrt{2e}\simeq 0.4289$), the scaled self-dual GKP codes achieve the lower bound in Eq. (2) exactly. However in all other cases, the achievable rate of the scaled self-dual GKP codes falls short of the lower bound of $Q(\mathcal{N}_{\sigma})$. Thus in what follows, we aim to reduce the gap between the lower bound of $Q(\mathcal{N}_{\sigma})$ and the achievable rate of GKP codes. In particular we focus on the large $\sigma$ regime (e.g., $0.5\leq\sigma\leq 1/\sqrt{e}\simeq 0.6065$) and also discuss the prospect of achieving a non-zero rate in the range $1/\sqrt{e}\leq\sigma<1/\sqrt{2}$, thereby improving on the best-known lower bound, since this possibility is not ruled out by the upper bound in Eq. (2).
Achievable rates of concatenated multi-mode GKP codes–Here we investigate the performance of various concatenated multi-mode GKP codes such as the surface-square GKP codes, the $[[5,1,3]]$-hexagonal GKP code, and the color-hexagonal GKP codes. Importantly, we focus on codes that encode only one logical qubit over the entire $N$ modes for reasons to be made clear below. Moreover we implement a maximum likelihood decoder (MLD) Vuillot et al. (2019); Conrad et al. (2022) either exactly Bravyi et al. (2014) or approximately using a tensor-network decoder Bravyi et al. (2014); Chubb (2021) such that we can probe the ultimate limits on the error-correction capability of these concatenated GKP codes against displacement errors.
@@
In Fig. 1, we show the achievable rates of the surface-square, $[[5,1,3]]$-hexagonal, color-hexagonal GKP codes as a function of the noise standard deviation $\sigma$. Since all these instances of multi-mode GKP codes encode one logical qubit over $N$ modes, we compute the achievable rate of these codes using the hashing bound Bennett et al. (1996); Wilde (2013), i.e., $R=I_{c}(\mathcal{P}(\vec{p}),\hat{I}/2)/N$. Here, $\hat{I}/2$ is the maximally-mixed qubit state and $\mathcal{P}(\vec{p})$ is the single-qubit Pauli error channel obtained by applying an $N$-mode GKP encoding, $N$ independent copies of Gaussian random displacement channels, and a decoding operation. Here the division by $N$ reflects the fact that $N$ copies of $\mathcal{N}_{\sigma}$ are consumed to produce each $\mathcal{P}(\vec{p})$.
For the surface-square GKP codes, we use an exact MLD based on an idea similar to the one given in Ref. Bravyi et al. (2014) adapted to GKP codes. For the $[[5,1,3]]$-hexagonal GKP code, given its small code size, we perform brute-force calculations of the coset probabilities of all four logical single-qubit Pauli operators to perform MLD exactly. Lastly for the color-hexagonal GKP codes, we perform MLD approximately by using a variant of the tensor-network decoder in Ref. Chubb (2021); Chubb and Lapeyre (2022) adapted to GKP codes. The approximation is made when truncating the bond dimensions to a maximum allowed cutoff (64 in our work) during tensor-network contractions. See Ref. sup for more details on the decoding procedures we use and Ref. lat for the source code.
@@
The results in Fig. 1 thus clearly show the particular relevance of one-logical-qubit-into-many-oscillators codes when searching for GKP codes that might achieve (or even surpass) the best-known lower bound of $Q(\mathcal{N}_{\sigma})$ in the large $\sigma$ regime, i.e., $\sigma\sim 1/\sqrt{e}\simeq 0.6065$. In particular, our results provide a non-trivial demonstration that GKP codes can achieve non-zero quantum state transmission rates for a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ at larger values of $\sigma$ than previously demonstrated Gottesman et al. (2001); Harrington and Preskill (2001); Harrington (2004) if a large number of modes (e.g., $N=1521$ for the $d=39$ surface-square GKP code) is used to encode just one logical qubit.
The largest multi-mode GKP code we could analyze numerically is the $d=39$ surface-square GKP code which achieves a non-zero quantum state transmission rate at a value of $\sigma$ as large as $\sigma=0.598$ (see Ref. sup ). However, higher performance may be achieved with an even larger number of modes or with another code family outperforming the surface-square GKP code family. In the rest of the paper, we provide additional numerical data to shed further light on the prospect of achieving (or surpassing) the lower bound of $Q(\mathcal{N}_{\sigma})$ with multi-mode GKP codes.
@@
Exact MLD for the surface-square GKP codes–We begin by illustrating the importance of using an exact MLD as opposed to a suboptimal decoder when exploring the quantum capacity of a Gaussian random displacement channel with GKP codes. In Fig. 2(a), we compare the performance of the surface-square GKP code with two different decoders, i.e., the closest-point decoder (CPD; dashed line) Lin et al. (2023) and the exact MLD (solid line) based on Ref. Bravyi et al. (2014). Since CPD converges to MLD only in the $\sigma\rightarrow 0$ limit, MLD outperforms CPD in the large $\sigma$ regime where $Q(\mathcal{N}_{\sigma})$ nearly vanishes. Interestingly when the exact MLD is used, the error threshold of the surface-square GKP code is very close to $\sigma=1/\sqrt{e}\simeq 0.6065$ at which the best-known lower bound of $Q(\mathcal{N}_{\sigma})$ vanishes.
@@
To get further insights into the asymptotic behavior of the surface-square GKP code in the $d\rightarrow\infty$ limit, we show in Fig. 2(b) the fidelity of the surface GKP code as a function of the code distance up to $d=39$ for three different values of $\sigma$, i.e., $\sigma\in[0.606,1/\sqrt{e}\simeq 0.6065,0.607]$. Remarkably, at exactly $\sigma=1/\sqrt{e}\simeq 0.6065$, the fidelity plateaus within the statistical variation (caused by a finite number of Monte Carlo samples) as the code distance increases. Moreover at a slightly smaller (or larger) value of $\sigma$, the fidelity increases (or decreases) as the code distance $d$ increases. These numerical observations suggest that for any $\sigma<1/\sqrt{e}$, the surface-square GKP code may be below the error threshold of $\mathcal{N}_{\sigma}$ and may consequently achieve a non-zero quantum state transmission rate with a sufficiently large code distance $d$. Note that our method of determining the threshold of the surface-square GKP code is more rigorous compared to the ones used in related prior works Fukui et al. (2017, 2018); Hänggli et al. (2020); Zhang et al. (2023), which studied several families of GKP codes against $\mathcal{N}_{\sigma}$ near $\sigma=1/\sqrt{e}\simeq 0.6065$ sup .

### v2

@@ I Introduction
Quantum communication [1, 2] and quantum computing [3, 4, 5] are two intricately related fields within quantum information science, aiming to leverage the principles of quantum mechanics to revolutionize data transmission and information processing. One common challenge in these areas is the need to protect the encoded quantum information from noise in quantum systems. Quantum error correction (QEC) thus plays a pivotal role in both fields by providing a mean to systematically suppress the effects of noise. Owing to this close relationship, the concepts and methodologies developed in one area can often be applied to another. For example, noisy processes in quantum computers can be modeled as noisy quantum channels for transmitting quantum information. This analogy is particularly fruitful when it is complemented by a general framework of quantum communication theory. In particular, the latter allows one to study maximum achievable data transmission rates of a noisy quantum channel through various entropic quantities [6, 7, 8, 9, 10], providing fundamental bounds on the capability of a QEC code for correcting errors in the noisy quantum channel.
@@
Bosonic Gaussian channels [11] have been extensively studied in the field of quantum communication theory [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] due to their relevance to realistic optical and microwave quantum communication channels. Moreover dominant errors in bosonic modes in various quantum computing platforms (e.g., 2D resonators [27] and 3D cavities [28, 29, 30] in superconducting systems, motional modes in trapped-ion systems [31, 32], and light modes in optical systems [33]) are well described by bosonic Gaussian channels. Thus, numerous bosonic QEC codes [34, 35, 13, 36, 37, 38] have been proposed to correct for errors in these bosonic Gaussian channels.
@@
In this work, we explore the interface between quantum communication theory and quantum error correction through bosonic Gaussian channels and bosonic QEC. In particular, we consider the problem of determining the quantum capacity [6, 7, 8, 9] $Q(\mathcal{N}_{\sigma})$ of a Gaussian random displacement channel $\mathcal{N}_{\sigma}$, i.e., its maximum achievable quantum state transmission rate per channel use under an optimal QEC strategy. The quantum capacities of some special cases of bosonic Gaussian channels, such as the pure-loss and pure-amplification channels, are well understood analytically [12, 17, 19, 20] thanks to their degradability or anti-degradability property [16]. However generic bosonic Gaussian channels including the Gaussian random displacement channel do not possess this special property. Thus only lower [12, 23] and upper [12, 19, 20, 21, 22] bounds on their quantum capacity are known.
@@
Here we attempt to make progress on this long-standing open problem by investigating the performance of multi-mode Gottesman-Kitaev-Preskill (GKP) codes [14, 15, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] against errors caused by a Gaussian random displacement channel $\mathcal{N}_{\sigma}$. Specifically, we aim to find a tighter lower bound on the quantum capacity of the Gaussian random displacement channel by searching for explicit QEC schemes which achieve higher quantum state transmission rates than the best-known lower bound on $Q(\mathcal{N}_{\sigma})$. Our search is based on multi-mode GKP codes and their maximum likelihood decoding.
II Gaussian random displacement channel and scaled self-dual GKP codes
@@
In the Heisenberg picture, a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ transforms the position and momentum operators $\hat{q}$ and $\hat{p}$ of a bosonic mode into $\hat{q}+\xi_{q}$ and $\hat{p}+\xi_{p}$, respectively. Here, $\xi_{q}$ and $\xi_{p}$ are independent classical noise variables following a Gaussian distribution with vanishing mean and a variance of $\sigma^{2}$. Note that a Gaussian random displacement channel is also referred to as an additive Gaussian noise channel in the literature.
@@
Here, $I_{c}(\mathcal{N},\hat{\sigma})\equiv S(\mathcal{N}(\hat{\sigma}))-S(\mathcal{N}^{c}(\hat{\sigma}))$ is the coherent information of the channel $\mathcal{N}$ with respect to an input state $\hat{\sigma}$, $S(\hat{\sigma})\equiv-\text{Tr}[\hat{\sigma}\log_{2}\hat{\sigma}]$ is the quantum von Neumann entropy, and $\mathcal{N}^{c}$ is the complementary channel of $\mathcal{N}$ [6, 7, 8, 9]. However in practice, evaluating this entropic quantity is difficult due to the maximization over all possible input states $\hat{\rho}$ as well as the need to consider an asymptotic limit of $N\rightarrow\infty$.
@@
where the lower bound is obtained by computing its one-shot coherent information (i.e., a similar expression as in Eq. (1) but with $N=1$, see App. H for a derivation) [12] and the upper bound is derived using a data-processing inequality [22]. From the lower bound and the definition of the quantum capacity, it follows that there must exist a QEC scheme that achieves a quantum state transmission rate of $\log_{2}(\frac{1}{e\sigma^{2}})$ for the Gaussian random displacement channel $\mathcal{N}_{\sigma}$ with $\sigma\leq 1/\sqrt{e}\simeq 0.6065$. Earlier works showed that scaled self-dual GKP codes can achieve the rate of $\log_{2}(\lfloor\frac{1}{e\sigma^{2}}\rfloor)$ [14, 15]. Notably, the floor function here is due to the fact that scaled self-dual GKP codes are restricted to cases where every single one of the $N$ normal modes encodes the same number (a positive integer $\lambda$) of logical states, thereby encoding $\lambda^{N}$ logical states or $N\log_{2}\lambda$ logical qubits overall. Thus when $\frac{1}{e\sigma^{2}}$ is given by a positive integer (e.g., $\frac{1}{e\sigma^{2}}=2\leftrightarrow\sigma=1/\sqrt{2e}\simeq 0.4289$), the scaled self-dual GKP codes achieve the lower bound in Eq. (2) exactly. However in all other cases, the achievable rate of the scaled self-dual GKP codes falls short of the lower bound of $Q(\mathcal{N}_{\sigma})$. Thus in what follows, we aim to reduce the gap between the lower bound of $Q(\mathcal{N}_{\sigma})$ and the achievable rate of GKP codes. In particular we focus on the large $\sigma$ regime (e.g., $0.5\leq\sigma\leq 1/\sqrt{e}\simeq 0.6065$) and also discuss the prospect of achieving a non-zero rate in the range $1/\sqrt{e}\leq\sigma<1/\sqrt{2}$, thereby improving on the best-known lower bound, since this possibility is not ruled out by the upper bound in Eq. (2).
III Achievable rates of concatenated multi-mode GKP codes
Here we investigate the performance of various concatenated multi-mode GKP codes such as the surface-square GKP codes, the $[[5,1,3]]$-hexagonal GKP code, and the color-hexagonal GKP codes. Importantly, we focus on codes that encode only one logical qubit over the entire $N$ modes for reasons to be made clear below. Moreover we implement a maximum likelihood decoder (MLD) [42, 46] either exactly [49] or approximately using a tensor-network decoder [49, 50] such that we can probe the ultimate limits on the error-correction capability of these concatenated GKP codes against displacement errors. We review MLD in the context of GKP codes in App. B.
@@ III Achievable rates of concatenated multi-mode GKP codes
In Fig. 1, we show the achievable rates of the surface-square, $[[5,1,3]]$-hexagonal, color-hexagonal GKP codes as a function of the noise standard deviation $\sigma$. Since all these instances of multi-mode GKP codes encode one logical qubit over $N$ modes, we compute the achievable rate of these codes using the hashing bound [51, 10], i.e.,
$R=\frac{1}{N}I_{c}\left(\mathcal{P}_{\vec{p}},\frac{1}{2}\hat{I}\right)=\frac{1}{N}\left(1+\sum_{\mu=I,X,Y,Z}p_{\mu}\log_{2}p_{\mu}\right).$ (3)
Here, $\hat{I}/2$ is the maximally-mixed qubit state and $\vec{p}=(p_{I},p_{X},p_{Y},p_{Z})$ are parameters of the single-qubit Pauli error channel $\mathcal{P}_{\vec{p}}$. The channel is obtained by applying an $N$-mode GKP encoding, $N$ independent copies of Gaussian random displacement channels, and a decoding operation. The parenthesis in Eq. (3) is the hashing bound of Pauli channels (see App. G for the derivation), which is divided by $N$ to reflect the fact that $N$ copies of $\mathcal{N}_{\sigma}$ are consumed to produce each $\mathcal{P}_{\vec{p}}$.
For the surface-square GKP codes, we use an exact MLD based on an idea similar to the one given in Ref. [49] adapted to GKP codes. For the $[[5,1,3]]$-hexagonal GKP code, given its small code size, we perform brute-force calculations of the coset probabilities of all four logical single-qubit Pauli operators to perform MLD exactly (see App. E for more details). Lastly for the color-hexagonal GKP codes, we perform MLD approximately by using a variant of the tensor-network decoder in Ref. [50, 52] adapted to GKP codes. The approximation is made when truncating the bond dimensions to a maximum allowed cutoff (64 in our work) during tensor-network contractions. See App. D for more details on the decoding procedures we use and Ref. [53] for the source code.
@@
The results in Fig. 1 thus clearly show the particular relevance of one-logical-qubit-into-many-oscillators codes when searching for GKP codes that might achieve (or even surpass) the best-known lower bound of $Q(\mathcal{N}_{\sigma})$ in the large $\sigma$ regime, i.e., $\sigma\sim 1/\sqrt{e}\simeq 0.6065$. In particular, our results provide a non-trivial demonstration that GKP codes can achieve non-zero quantum state transmission rates for a Gaussian random displacement channel $\mathcal{N}_{\sigma}$ at larger values of $\sigma$ than previously demonstrated [13, 14, 15] if a large number of modes (e.g., $N=1521$ for the $d=39$ surface-square GKP code) is used to encode just one logical qubit.
The largest multi-mode GKP code we could analyze numerically is the $d=39$ surface-square GKP code which achieves a non-zero quantum state transmission rate at a value of $\sigma$ as large as $\sigma=0.598$ (see App. A). However, higher performance may be achieved with an even larger number of modes or with another code family outperforming the surface-square GKP code family. In the rest of the paper, we provide additional numerical data to shed further light on the prospect of achieving (or surpassing) the lower bound of $Q(\mathcal{N}_{\sigma})$ with multi-mode GKP codes.
@@ IV Exact MLD for the surface-square GKP codes
We begin by illustrating the importance of using an exact MLD as opposed to a suboptimal decoder when exploring the quantum capacity of a Gaussian random displacement channel with GKP codes. In Fig. 2(a), we compare the performance of the surface-square GKP code with two different decoders, i.e., the closest-point decoder (CPD; dashed line) [48] and the exact MLD (solid line) based on Ref. [49]. More details on the exact MLD can be found in App. C. Since CPD converges to MLD only in the $\sigma\rightarrow 0$ limit, MLD outperforms CPD in the large $\sigma$ regime where $Q(\mathcal{N}_{\sigma})$ nearly vanishes. Interestingly when the exact MLD is used, the error threshold of the surface-square GKP code is very close to $\sigma=1/\sqrt{e}\simeq 0.6065$ at which the best-known lower bound of $Q(\mathcal{N}_{\sigma})$ vanishes.
@@
To get further insights into the asymptotic behavior of the surface-square GKP code in the $d\rightarrow\infty$ limit, we show in Fig. 2(b) the fidelity of the surface GKP code as a function of the code distance up to $d=39$ for three different values of $\sigma$, i.e., $\sigma\in[0.606,1/\sqrt{e}\simeq 0.6065,0.607]$. Remarkably, at exactly $\sigma=1/\sqrt{e}\simeq 0.6065$, the fidelity plateaus within the statistical variation (caused by a finite number of Monte Carlo samples) as the code distance increases. Moreover at a slightly smaller (or larger) value of $\sigma$, the fidelity increases (or decreases) as the code distance $d$ increases. These numerical observations suggest that for any $\sigma<1/\sqrt{e}$, the surface-square GKP code may be below the error threshold of $\mathcal{N}_{\sigma}$ and may consequently achieve a non-zero quantum state transmission rate with a sufficiently large code distance $d$. Note that our method of determining the threshold of the surface-square GKP code is more rigorous compared to the ones used in related prior works [39, 40, 44, 54], which studied several families of GKP codes against $\mathcal{N}_{\sigma}$ near $\sigma=1/\sqrt{e}\simeq 0.6065$ (See App. F for further discussions).
V Approximate MLD for color-hexagonal GKP codes using a tensor-network decoder